OpenStudy (anonymous):

A store mixes Brazilian coffee worth $10 per kilogram and Turkish coffee worth $12 per kilogram. The mixture is to sell for $11 per kilogram. Find how much of each should be used to make a 396-kilogram mixture

OpenStudy (anonymous):

well, if it is simply asking for how much BEANS, then its quite easy, just divide 396 by two because the mixture uses 1 kilogram of each. They tried to trick you, I had to look at it for a while.

OpenStudy (anonymous):

@HyperChemist Suppose the cost of Brazillian and Turkish coffee was $10.4 and $17 per kg respectively. What would be the weight of each coffee to make a 396 kg mixture?

OpenStudy (anonymous):

@lattrece Let B and T be the weight of the cost of Brazillian and Turkish coffee per kg respectively. Solve the following equations for B and T: { 10 B + 12 T = 11 ( B + T ), B + T = 396} Both B and T = 198 kg.